OpenStudy (zeig_101):

Help please! Does the series converge or diverge? If it converges, what is the sum? (infinity) sigma -4(-(1/2))^(n-1) n-1

OpenStudy (anonymous):

it's a geometric series with the common ratio = -1/2. what does that tell you?

OpenStudy (zeig_101):

each number in the sequence is *-.5 from the last

OpenStudy (anonymous):

ok, so how do you know if the series converges or diverges?

OpenStudy (zeig_101):

I don't recall.

OpenStudy (anonymous):

if |r| < 1, then the series converges. Otherwise, diverges

OpenStudy (zeig_101):

So it converges?

OpenStudy (anonymous):

is |-1/2| < 1?

OpenStudy (zeig_101):

Yes, but it's to the power of n-1, and multiplied by -4

OpenStudy (anonymous):

-4 is irrelevant. it's the common ratio that determines whether a geometric series converges or diverges

OpenStudy (zeig_101):

okay, so it converges because -0.5<1

OpenStudy (anonymous):

so, it converges because |-1/2| < 1

OpenStudy (zeig_101):

How do I find the sum if it uses infinity, though?

OpenStudy (anonymous):

do you know the formula?

OpenStudy (zeig_101):

I think I do.

OpenStudy (anonymous):

ok what is it?

OpenStudy (zeig_101):

x(n)=ar^(n-1)

OpenStudy (anonymous):

no, it's sum = (first term) / (1 - common ratio)

OpenStudy (zeig_101):

X=n/1=-0.5

OpenStudy (zeig_101):

ok

OpenStudy (anonymous):

sum = -4/(1 - (-1/2)) do them math

OpenStudy (zeig_101):

-(8/3)?

OpenStudy (anonymous):

yea

OpenStudy (zeig_101):

Is that the sum?

OpenStudy (anonymous):

yea

OpenStudy (zeig_101):

Alright, thanks!

OpenStudy (anonymous):

np